OpenStudy (anonymous):

Which algebraic expression would not be equivalent if the number and the variable were reversed? A. n + 8 B. m – 5 C. k x 8 D. 4 + x

OpenStudy (anonymous):

addition and multiplication are commutative(means the equation remains same even when reversed), subtraction and division are not. So what will the answer be?

OpenStudy (anonymous):

m-5

OpenStudy (anonymous):

thats right

OpenStudy (midhun.madhu1987):

Commutative Property is NOT satisfied for Subtraction and Division... Example: x + 9 = 9 + x x * 9 = 9 * x But x - 9 Not equal to 9 - x